OpenStudy (anonymous):

explain how decreasing the coefficient of x affects the graph of the equation y=|x|

OpenStudy (campbell_st):

well you are really graphing 2 lines y = x and y = -x they intersect at the origin so using your knowledge of the slope of a straight line what would you say if you are comparing the lines above with y = 1/2 x and y = -1/2 x